As I explained in class, Treasury Inflation Protected Security(TIPS) issued by the U.S. Governmentoffersprotection from decayin the real purchasing power of an investment.Please estimatepromised YTM on a1-yr T-Noteunder assumption that 1-yr TIPS promises to pay 2.0% and investors anticipate 3.0% rate of inflation over the same time period.

A) 0.0506%B) 1.0506%C) 5.00%D) 5.06%E) None of the above
Answer #1

D)5.06%

An inflation adjusted bond will adjust the returns based on inflation rate.

Since in the above question maturity time is 1 year we can directly calculate yield.

Let the par value be $1

If the inflation rate is 3% then the par value will be adjusted to

Par value=1+.03×1=$1.03

Given rate of return=2% on the new par value.

Therefore value of bond after maturity=1.03+.02×1.03

=$1.0506

Therefore yield=net return=(final value-initial value)÷initial value=(1.0506-1)÷1=.0506

=5.06%

Thanks for the question. Keep learning.